What is the slope-intercept form of the equation of the line shown in the graph? Coordinates are (3,0) and (0, -4)???? Can someone help me?
Answer by Abbey(339)

find the slope by taking the difference of the y values/difference of x values:
(y-y1)/(x-x1)
(0-(-4))/(3-0) = 4/3
find the intercept by plugging one of these points and the slope into the equation y=mx+b
0=(4/3)*3+b
0=4+b
b=-4
then the formula is
y=mx+b
y=(4/3)x -4